(298k) 9. What was the volume at STP of a sample of gas that occupies 16.5L at 352°C and 0.275atm?
Answer #1

P1=0.275 atm, V1=16.5L ,T1=3520C =352+273=625K

at STP P2=1 atm, T2=273K ,V2=?

as P1*V1/T1 = P2*V2/T2

V2= (P1*V1*T2)/(T1*P2) =(0.275atm*16.5*273K)/(625K*1atm) =1.98L

So answer is 1.98L

thank you